OpenStudy (anonymous):

choosing one of three science courses one of five mathmatics courses one of two english courses and one of four soft drinks

OpenStudy (anonymous):

huh? what is the relation btwn science courses and soft drinks?

OpenStudy (anonymous):

is that exactly what the question says?

OpenStudy (anonymous):

i mean four social studiescourses

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

counting princple.

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

\[3\times 5\times 2\times 4\]

OpenStudy (anonymous):

thanks satellite

OpenStudy (anonymous):

yw